OpenStudy (shadow13):

1. The length of a rectangle is 5 mm longer than its width. Its perimeter is more than 30 mm. Let w equal the width of the rectangle. a. Write an expression for the length in terms of the width. b. Use expressions for the length and width to write an inequality for the perimeter, on the basis of the given information. c. Solve the inequality, clearly indicating the width of the rectangle.

OpenStudy (shadow13):

@kimberlyevens

OpenStudy (kimberlyevens):

Given: width = w length = w + 5mm perimeter > 30 mm Perimeter = 2 ( l + w) 2 ((w+5) + w) > 30 2 (2w + 5) > 30 4w + 10 > 30 - 10 -10 4w > 20 ÷4 ÷4 w > 5 width should be greater than 5 so that the perimeter will be greater than 30. To check: assuming width is 6 mm w = 6 mm l = w + 5 mm ⇒ 6 mm + 5 mm = 11 mm Perimeter = 2 (l + w) ⇒ 2 (11 + 6) = 22 + 12 = 34 mm 33 mm > 30 mm

OpenStudy (shadow13):

@imqwerty

OpenStudy (shadow13):

is that right?

OpenStudy (kimberlyevens):

I believe so I do math problems like this all the time

OpenStudy (shadow13):

okay thanks ill put it in my words or u want me to just copy and paste it?

OpenStudy (kimberlyevens):

what is it for

OpenStudy (shadow13):

its for an test

OpenStudy (kimberlyevens):

like a quiz or a document

OpenStudy (shadow13):

document

OpenStudy (kimberlyevens):

I would rewrite it then because if you copy and past that's plagiarizing and if your teacher finds out you will get a bad mark on your record

OpenStudy (shadow13):

kk

OpenStudy (kimberlyevens):

good luck:)

OpenStudy (danjs):

i would try to understand the problem and how it was solved. Then do it by yourself on the exam without help.